A 1.5x10 3C charge needs to move 0.25m from certain point. Calculate it s electric potential voltage.
1
Expert's answer
2021-10-15T10:22:32-0400

The electric potential

"V=k\frac{q}{r}=9*10^9*\frac{1.5*10^{-3}}{0.25}=5.4*10^7\:\rm V"